OpenStudy (anonymous):

Mrs. Fowler has 28 students in her class, and 16 of them are boys. What is the ratio of girls to boys? A. 4 : 3 B. 3 : 4 C. 4 : 11 D. 3 : 11

OpenStudy (pinklion23):

hey im back

OpenStudy (igreen):

Girls = 28 - 16 = 12 Boys = 16 Ratio: 12:16 Simplify by dividing 4 to both numbers. 12 / 4 = ? 16 / 4 = ?

OpenStudy (anonymous):

what @pinklion23

OpenStudy (pinklion23):

yeah iGreen is right and he explained it very simply

OpenStudy (pinklion23):

do you get it?

OpenStudy (anonymous):

3 4

OpenStudy (pinklion23):

YEAH! you did it good job

OpenStudy (firejay5):

@Sworewolf The correct answer is B. 3:4, because 12/4 = 3 = girls and 16/4 = 4 = guy/boys